OpenStudy (anonymous):

solve the system of equation, whats the solution. y=-x+21 y=-2x+30

OpenStudy (anonymous):

y=-x+21 move y to the right side and x to the left side x=21-y then put this in the second equation y=-2(21-y)+30 y=-42+2y+30 y-2y=-42+30 -y=-12 y=12 x=21-y so x=21-12 x=9

OpenStudy (anonymous):

tank you c:

OpenStudy (anonymous):

u r welcome

OpenStudy (anonymous):

(: